### v4

# An exact analytical solution for Dicke superradiance

Abstract
We revisit the Dicke superradiance problem, namely finding the time evolution of the density operator for an initially inverted ensemble of $N$ identical two-level systems (spins), undergoing collective spontaneous decay. We derive a compactly written, easily computable and fully analytical solution, valid for any time during the evolution and for any $N$.
Dicke superradiance refers to the quick burst of radiation from an initially fully inverted ensemble of $N$ collectively decaying two-level quantum systems, or spins, with individual rate $\Gamma$. This feature has been introduced by Dicke in 1954 [1] and first analytically solved by Lee in 1977, starting with the full excitation case [2] and generalizing to any arbitrary initial state [3]. Other analytical treatments followed, for example described in Refs. [4, 5].
To be more precise, the Dicke superradiance problem involves finding the density operator $\rho$ which satisfies the following master equation in Lindblad form (in an interaction picture where the trivial action of the free Hamiltonian is removed):
$\dot{\rho}(t)=\mathcal{L}[\rho]=\Gamma\Bigg{[}S\rho S^{\dagger}-\frac{1}{2}\Big{(}S^{\dagger}S\rho-\rho S^{\dagger}S\Big{)}\Bigg{]}.$ (1)
The ladder operators in the symmetric subspace spanned by $N+1$ collective Dicke states |$\ket{m}$ indexed by $m=0,1,...,N$ have the following action $S|m\rangle=\sqrt{h_{m}}|m-1\rangle$ and $S^{\dagger}|m\rangle=\sqrt{h_{m+1}}|m+1\rangle$. The coefficients $h_{m}=m\bar{m}$ with $\bar{m}=(N+1)-m$ are double degenerate (for even $N$) as $h_{m}=h_{\bar{m}}$. For odd $N$, there is a point of no degeneracy at $(N+1)/2$.
The solution proposed in Refs. [2, 3] has been mostly overlooked during the last decades; presumably the convoluted analytical expressions renders this solution computationally demanding, thus requiring resources similar to the direct simulation of the master equation, which is a trivial task for particles numbers of the order $N=10^{3}$ spins. The focus instead shifted more recently to deriving more informative solutions in the mesoscopic (or thermodynamic) limit of large $N$, such as, for example, in Refs. [6, 7, 8, 9, 10, 11]. A comprehensive early review of the problem and its implications can be found in Ref. [12].
While the original formulation of the Dicke superradiance model is impossible to realize, as it involves $N$ atoms spatially positioned at same point, a straightfoward realization can be achieved in cavity quantum electrodynamics, such as for example with atoms trapped in the standing wave field of a lossy optical cavity (with coupling and loss rates $g$ and $\kappa$, respectively). The cavity opens a decay channel with a collective collapse operator $S$ (after adiabatic elimination of the cavity mode) and a collapse rate $\Gamma=g^{2}/\kappa$. Assuming negligible independent spontaneous emission, the problem is described by Eq. (1). Such collective cavity superradiance effect has been recently implemented to advance superradiant lasers. Moreover, waveguide and circuit quantum electrodynamics provide alternative platforms to realize Dicke superradiance, where excited (artificial) atoms decay collectively into a nearby microwave or optical resonator mode.
Solution – We provide here an exact analytical solution, expressed in a compact form and numerically tractable. The diagonal elements of the total density operator $\rho(t)=\sum_{m=0}^{N}\rho_{m}(t)|m\rangle\langle m|$ assume the following solution
$\rho_{m}(t)=\sum_{j=j_{m}}^{N}\text{Res}[f_{m}(z,t)]\Bigg{|}_{z=h_{j}},$ (2)
where $j_{m}=\mathrm{max}\{m,N/2+1\}$, for even $N$ and $j_{m}=\mathrm{max}\{m,(N+1)/2\}$, for odd $N$. The terms in the sum are the residues of a function $f_{m}(z,t)$ (which we list below) around poles situated at $h_{j}$, of single or double multiplicity. More explicitly, for even $N$ (while the odd case is presented in the Appendix)
$\rho_{m}(t)=\sum_{j=m}^{N}\Bigg{[}f_{m}(z,t)(z-h_{j})\Bigg{]}_{z=h_{j}}\theta_{m>\frac{N}{2}}+\Bigg{\{}\sum_{j=\bar{m}+1}^{N}\Bigg{[}f_{m}(z,t)(z-h_{j})\Bigg{]}_{z=h_{j}}+\sum_{j=N/2+1}^{\bar{m}}\frac{d}{dz}\Bigg{[}f_{m}(z,t)(z-h_{j})^{2}\Bigg{]}_{z=h_{j}}\Bigg{\}}\theta_{m\leq\frac{N}{2}}.$ (3)
As illustrated in Fig. 1, for points above the equator ($m>N/2$), the poles are non-degenerate. The Heaviside $\theta_{m>N/2}$ function is defined as $1$ for any $m\geq N/2+1$ and elsewhere zero. Below the equator we use $\theta_{m\leq N/2}$ being $1$ for $m$ between $1$ and $N/2$ and zero elsewhere. Some of the poles are second order given that in the interval from $m$ to $\bar{m}$ there exist a number of degenerate values $h_{j}=h_{\bar{j}}$. The generating function on which the residues are to be computed is
$f_{m}(z,t)=(-1)^{m}\frac{h_{N}...h_{m+1}}{(z-h_{N})\ldots(z-h_{m})}e^{-z\Gamma t}.$ (4)
The solution above is easily numerically computable; however, we will in the following express it more clearly in terms of linear combinations of the $N$ eigenvalues of the Lindblad superoperator, i.e. the set of collective exponential decay channels characterized by the values $e^{-h_{m}\Gamma t}$, for $m=1,\ldots N$.
Figure 1: Left: Contributions to the solution in Eq. (3) above the equator adding only simple poles marked in blue. Right: Separation of simple pole contributions in blue and double poles contributions ($h_{j}=h_{\bar{j}}$) marked in red, for even $N$.
Derivation – Let us now describe the analytical procedure to reach the solution presented above. The $N+1$ diagonal elements of the density matrix can be written formally as
$\rho_{m}(t)=\langle m|e^{\mathcal{L}t}\rho(0)|m\rangle=\sum_{j=0}^{\infty}\frac{(\Gamma t)^{j}}{j!}\rho_{m}^{(j)},$ (5)
where we expressed each element as a time series with coefficients $\rho_{m}^{(j)}=\langle m|\mathcal{L}^{j}[\rho(0)]|m\rangle=\langle m|\mathcal{L}[\mathcal{L}[\cdots\mathcal{L}[\rho(0)]]]|m\rangle$ indicating that the Lindblad superoperator is applied $j$ times to the initially fully excited state $\rho(0)=|N\rangle\langle N|$ of the system. We start by noticing the rule that
$\mathcal{L}[|m\rangle\langle m|]=-h_{m}|m\rangle\langle m|+h_{m}|m-1\rangle\langle m-1|,$ (6)
indicating that the collective decay moves the system either horizontally within the same Dicke state at a rate $h_{m}$ or diagonally to the lower Dicke state with $m-1$ at a rate $h_{m}$. Of course, at $m=0$, at the bottom of the ladder, we have $\mathcal{L}[|0\rangle\langle 0|]=0$. We can then find a general recursion rule
$\rho_{m-1}^{(j+1)}=-h_{m-1}\rho^{(j)}_{m-1}+h_{m}\rho^{(j)}_{m},$ (7)
which shows that any order $j+1$ in the time power series can be computed from the previous order $j$ with knowledge of all the terms $\rho_{m}^{(j)}$. Starting with all emitters excited at time $t=0$ means that $\rho_{N}^{(0)}=1$, which allows one to compute $\rho_{N}^{(1)}=-h_{N}$ and so on generally leading to $\rho_{N}^{(j)}=(-h_{N})^{j}$. The summation of the infinite sum leads to the expected result $\rho_{N}(t)=\mathrm{exp}(-h_{N}\Gamma t)=\mathrm{exp}(-N\Gamma t)$.
Solving the recursive equation (see Appendix) indicates a general form for the solution (of limited direct validity, as explained below)
$\rho_{m}(t)=\sum_{j=m;j\neq m}^{N}(-1)^{m}\frac{h_{N}...h_{m+1}}{(h_{j}-h_{N})\ldots(h_{j}-h_{m})}e^{-h_{j}\Gamma t}.$ (8)
The solution above proves correct for any point $m\geq N/2+1$ but it fails below the equator. However, we start with the observation that the expression above is nothing else but the sum over first order residues of the extended function $f_{m}(z,t)=\mathcal{P}_{m}(z)e^{-z\Gamma t}$. We introduce the notation
$\Bigg{[}f(z,t)(z-h_{j})\Bigg{]}_{z=h_{j}}=\mathcal{P}_{mj}e^{-h_{j}\Gamma t}.$ (9)
where the elements
$\mathcal{P}_{mj}=(-1)^{m}\frac{h_{N}\ldots h_{m+1}}{(h_{j}-h_{N})\ldots 1\ldots(h_{j}-h_{m})},$ (10)
are obtained by removing the term ($h_{j}-h_{j}$) (replacing it with $1$) in the denominator of $\mathcal{P}_{m}(h_{j})$.
Below the equator, we identify instead double poles in the sums running over indices with $m\leq N/2$, and apply the residue procedure taking into account this aspect. This is indicated in Fig. 1 where the index $j$ in the sum runs over simple poles between $\bar{m}+1$ and $N$ and over doubly degenerate poles in the rest of the sum. The residues of the double poles are given by
$\frac{d}{dz}\Bigg{[}f(z,t)(z-h_{j})^{2}\Bigg{]}\Bigg{|}_{z=h_{j}}=\mathcal{R}_{mj}(t)e^{-h_{j}\Gamma t}.$ (11)
Notice that the terms multiplying the exponentials are now matrix elements with implicit time dependence
$\mathcal{R}_{mj}(t)=-\Gamma t\tilde{\mathcal{P}}_{mj}-\frac{d\tilde{\mathcal{P}}_{m}(z)}{dz}\Bigg{|}_{z=h_{j}}.$ (12)
The function $\tilde{\mathcal{P}}_{m}(z)$ is derived from $\mathcal{P}_{m}(z)$ by removing two terms at $j$ and $\bar{j}$ in the denominator. The matrix elements $\tilde{\mathcal{P}}_{mj}$ similarly are obtained by removing the two terms in the denominator and then evaluating it at $z=h_{j}$. Explicitly one can then write
$\tilde{\mathcal{P}}_{mj}=(-1)^{m}\frac{h_{N}\ldots h_{m+1}}{(h_{j}-h_{N})\ldots 1\ldots 1\ldots(h_{j}-h_{m})}.$ (13)
The analytical solution is easily simulated numerically for large values of $N$, well above the values tractable by direct simulation of the master equation. An example is shown in Fig. 2 where the distribution of the Dicke state populations of the density operator at the time of peak emission shown for $N$ increasing from $10^{2}$ to $10^{4}$.
Figure 2: Distribution of Dicke states at the peak emission time, computed from the analytical expression in Eq. (3). The distrbution shows a more pronounced peak as the particle number is increased. The emission rate is computed as $I=\Gamma\langle S^{\dagger}S\rangle(t)$ and it peaks around $t_{\text{peak}}\approx\log{N}/(N\Gamma)$. We observe, that the distributions peak at $m=39$, $m=362$ and $m=3414$ respectively.
Matrix formulation – We can express the solution also in a more convenient form by ordering the density operator diagonal elements in a vector of dimensions $N$ containing all states except for the ground state $m=0$, which can be obtained from the trace preservation constraint. This can be cast into a $N\times N$ lower triangular matrix as follows
$\begin{bmatrix}\rho_{N}(t)\\ \rho_{N-1}(t)\\ \vdots\\ \rho_{N/2+1}(t)\\ \hline\cr\rho_{N/2}(t)\\ \rho_{N/2-1}(t)\\ \vdots\\ \rho_{1}(t)\end{bmatrix}=\left[\begin{array}[]{c c c c|c c c c}\mathcal{P}_{N,N}&0&\ldots&0&0&0&\ldots&0\\ \mathcal{P}_{N-1,N}&\mathcal{P}_{N-1,N-1}&\ldots&0&0&0&\ldots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathcal{P}_{N/2+1,N}&\mathcal{P}_{N/2+1,N-1}&\ldots&\mathcal{P}_{N/2+1,N/2+1}&0&0&\ldots&0\\ \hline\cr\mathcal{P}_{N/2,N}&\mathcal{P}_{N/2,N-1}&\ldots&\mathcal{R}_{N/2,N/2+1}&\mathcal{R}_{N/2,N/2}&0&\ldots&0\\ \mathcal{P}_{N/2-1,N}&\mathcal{P}_{N/2-1,N-1}&\ldots&\mathcal{R}_{N/2-1,N/2+1}&\mathcal{R}_{N/2-1,N/2}&\mathcal{R}_{N/2-1,N/2-1}&\ldots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathcal{R}_{1,N}&\mathcal{R}_{1,N-1}&\ldots&\mathcal{R}_{1,N/2+1}&\mathcal{R}_{1,N/2}&\mathcal{R}_{1,N/2-1}&\ldots&\mathcal{R}_{1,1}\\ \end{array}\right]\begin{bmatrix}e^{-h_{N}\Gamma t}\\ e^{-h_{N-1}\Gamma t}\\ \vdots\\ e^{-h_{N/2+1}\Gamma t}\\ \hline\cr e^{-h_{N/2}\Gamma t}\\ e^{-h_{N/2-1}\Gamma t}\\ \vdots\\ e^{-h_{1}\Gamma t}\end{bmatrix}$ (14)
For numerical simulations, the elements of the matrix above can be written as $\mathcal{P}_{mj}\theta_{j\geq m}\theta_{j>\bar{m}}+\mathcal{R}_{mj}\theta_{j\geq m}\theta_{\bar{m}\geq j}$ where only the $\mathcal{R}_{mj}$ have an explicit time dependence.
Any initial state – Let us generalize our analytical solution to the case where the system is initialized in any Dicke state, denoted by the state vector |$\ket{m_{0}}$, such that $1\leq m_{0}\leq N$. To this end, we set $\rho_{m_{0}}=1$ and we derive the time evolution: the procedure is straightforward and it involves removing the terms pertaining to values between $m_{0}$ and $N$, thus obtaining
$\rho_{m}(t)=\sum_{j=j_{m}}^{m_{0}}\text{Res}[f_{m_{0},m}(z,t)]\Bigg{|}_{z=h_{j}}.$ (15)
Notice that the sum over the residues runs only to $m_{0}$ and the function is redefined as
$f_{m_{0},m}(z,t)=(-1)^{m_{0}+m}\frac{h_{m_{0}}...h_{m+1}}{(z-h_{m_{0}})\ldots(z-h_{m})}e^{-z\Gamma t}.$ (16)
Using this result, one can then with writing the time evolution of the full density matrix for any initial state written as a mixure of Dicke states $\rho(0)=\sum_{j=1}^{N}\rho_{j}(0)\ket{j}\bra{j}$ |.
Conclusions and outlook - We have provided a compact, fully analytical solution to the Dicke superradiance process, within the fully symmetric subspace of $N$ quantum emitters. The main observations have been i) the existence of simple recursive relations between Dicke states of index $m$ and $m-1$ (also previously observed in other works such as Refs. [11]), ii) the possibility to find a closed analytical solution for the recursive equation and iii) the identification of this solution as a sum of residues of a given simple function. This suggests that the solution could be expressed as a complex integral on a closed contour in the complex plane, which we will investigate in the future. We have also shown that the same simple formulation can be extended to the time evolution of any initial Dicke state as an initial state of the system.
The procedure followed here can be easily extended to provide fully analytical solutions to the case of the coherently driven Dicke superradiant system. Moreover, the analytical approach described here might provide ways to tackle the case of incoherently driven dissipative Dicke systems such is the case of superradiant lasing: in this case, $N$ quantum emitters are placed inside a lossy optical cavity or around optical fibers, where after the elimination of the cavity mode, the system behaves like a Dicke superradiant system with an external individual incoherent pump.
Acknowledgments - We acknowledge financial support from the Max Planck Society and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 429529648 – TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”). R.H. acknowledges funding by the Austrian Science Fund (FWF) 10.55776/W1259.